OpenStudy (anonymous):

find y' if y=sec2X/tan2X using the quotient rule within the chain rule

OpenStudy (asnaseer):

this might help you: http://www.alcyone.com/max/reference/maths/derivatives.html

OpenStudy (anonymous):

y=sec2X/tan2X , SEC 2X=1/COS 2X Y=1/(COS 2X)(TAN 2X) =1/(COS 2X)(SIN 2X /COS 2X) Y=1/SIN 2X=CSC 2X Y '= -2CSC 2X COT 2X ans............................. further simplification, using trig gives = 4cos 2x/(cos4x -1) ans